OpenStudy (anonymous):

PLEASE HELP: Find the solution to this exponential equations, correct to four decimal places. 10^(1-x) = 6^x

OpenStudy (saifoo.khan):

\[\large 1- x = \log_{10} 6 ^ x\] \[\large 1- x =x \log_{10} 6\]

OpenStudy (anonymous):

1 = xlog(6) +x 1 = x (log(6) +1) x = 1/(log(6)+1)

OpenStudy (anonymous):